Algebra: About 13 out of 20 homes have a personal computer.on a street with 60 homes, how many would you expect to have a personal computer?
SO, what I did was that i first lined up the proportion like this: 13/20 = x/60 but another person is saying it's like this: 13/x = 20/60
I am not sure which one is right...

Answer by jim_thompson5910(35256)

Both are the same because you can go from 13/20 = x/60 to 13/x = 20/60


Here's why

13/20 = x/60

13*60 = 20x ... Cross multiply

13*60/20 = x ... Divide both sides by 20

13/20 = x/60 ... Divide both sides by 60


So this work shows us that 13/20 = x/60 and 13/x = 20/60 are equivalent (which means the solutions to both are the same)

Let's solve the first one for x

13/20 = x/60

13*60 = 20x

780 = 20x

780/20 = x

39 = x

x = 39

Then let's solve the second equation for x

13/x = 20/60

13*60 = 20x

780 = 20x

780/20 = x

39 = x

x = 39


So either equation gives you the same correct value of x: that value is x = 39.

So on a street with 60 homes, you would expect 39 homes to have a personal computer.